# Linear-Cost Vecchia Approximation of Multivariate Normal Probabilities

### v2

@@ 1 Introduction
Current computational approaches for MVN probabilities are mostly based on the separation of variable (SOV) method (Genz, 1992). SOV applies a series of transformations such that the MVN probability can be written as an expectation of quantities with low variance, allowing an approximation of the expectation as the average of Monte Carlo (MC) samples. SOV consists of two stages: in the pre-processing stage, the Cholesky factor of the $n\times n$ MVN covariance matrix is computed in $\mathcal{O}(n^{3})$ time; in the MC sampling stage, $N$ samples are drawn, each in $\mathcal{O}(n^{2})$ time. The resulting SOV time complexity is $\mathcal{O}(n^{3}+Nn^{2})$. To improve upon the original SOV method, Gibson et al. (1994) introduced a univariate variable reordering technique used in the pre-processing stage that produces the Cholesky factor and further reduces the MC variance. Based on the series of integrand transformation used in the SOV method, Botev (2017) used importance sampling with minimax exponentially tilted (MET) normal proposal densities during the MC sampling stage to reduce variance. This results in substantially higher accuracy for tail MVN probabilities than SOV, but it requires iterative optimization of the proposal density in the pre-processing stage at a per-iteration complexity of $\mathcal{O}(n^{3})$; in practice, this optimization can be more expensive than the MC sampling for moderately large $n$ (e.g., $n>1{,}000$). Sampling from truncated MVN (TMVN) distributions is closely connected to approximating MVN probabilities (Botev, 2017). Specifically, the integrand in MET (or SOV) can be decomposed into two parts that can be viewed as a proposal density and a density ratio, which can be used for importance sampling from TMVN distributions at the same cost as estimating the corresponding MVN probability. MET is currently the state-of-the-art in terms of achieving the highest accuracy for estimating MVN probabilities and the highest acceptance rate for sampling TMVN distributions.
@@
For evaluating high-dimensional MVN probabilities with spatial covariance matrices, Genton et al. (2018) and Cao et al. (2021) used hierarchical-matrix approximations to reduce the complexity of pre-processing and MC sampling for SOV by a factor of $n^{1/2}$. Bolin & Lindgren (2015) re-parameterized the SOV method with the inverse Cholesky factor to take advantage of the sparse precision matrix under Gaussian Markov random fields. However, they found this approach to often be computationally inefficient, because the sparsity was not well preserved during Cholesky factorization. Other recent approaches for spatial MVN probabilities have used versions of the Vecchia approximation (Vecchia, 1988), which decomposes a multivariate density into a product of univariate conditional densities with small conditioning sets. Saha et al. (2022) applied a Vecchia-type approximation to reduce the complexity of computing the Cholesky factor in the preprocessing stage, but its per-sample complexity remains as $\mathcal{O}(n^{2})$. Nascimento & Shaby (2022) decomposed the $n$-dimensional MVN probability into a product of lower-dimensional MVN probability ratios through assuming conditional independence in the MVN cumulative distribution function (CDF); while this is based on a similar principle as the Vecchia approximation, it appears that the screening effect (i.e., far-away information contributes little after conditioning on nearby points) (Stein, 2002) is weaker for MVN CDFs than for densities, resulting in substantially increased approximation error for CDF approximations.
@@
Our proposed methods are also based on the Vecchia approximation, which makes a sparse-directed-acyclic-graph assumption. A Vecchia approximation of a MVN distribution results in a valid MVN distribution that has a covariance matrix with a sparse inverse Cholesky factor (Datta et al., 2016; Katzfuss & Guinness, 2021). Vecchia approximations can provide asymptotically exact inference in near-linear time for GPs with Matérn-type kernels up to boundary effects (Schäfer et al., 2021; Kang et al., 2024). Numerically, Vecchia approximations have been shown to be highly accurate even with small conditioning-set sizes in many settings, including spatial statistics and high-dimensional and multi-output GPs (e.g., Datta et al., 2016; Katzfuss & Guinness, 2021; Katzfuss et al., 2022; Schäfer et al., 2021; Cao et al., 2023; Kang & Katzfuss, 2023). Vecchia approximations can be applied to any covariance matrix, and we expect them to be most accurate if the covariance can be viewed as a stationary kernel with relatively small noise-to-signal ratio defined over some (potentially unknown) input space (Kang & Katzfuss, 2023).
@@
We propose scalable evaluation of MVN probabilities and sampling from TMVN distributions in linear time by applying suitable Vecchia approximations to the MET approach. Hence, we expect our methods to work well in the settings discussed in the previous paragraph. We show that the dominant computations for MET and SOV are equivalent to calculating the conditional means and variances of the integration variables. We reduce the conditioning sets in these calculations via the Vecchia approximation. We take full advantage of Vecchia’s sparse inverse Cholesky factor when computing the transformed MET integrand as well as its gradient and Hessian, which allows us the optimize the MET proposal density and sample from TMVN distributions in linear time, an $\mathcal{O}(n^{2})$ reduction compared to MET. We demonstrate theoretically that the Vecchia approximation error decreases exponentially with the number of nonzeros under certain regularity conditions, and we show numerically that the Vecchia error is often negligible relative to the Monte-Carlo error. In addition, both the pre-processing and sampling stages of our method are parallelizable. We show that our proposed method facilitates model estimation and posterior inference for partially censored MVN models that offer substantial information gain over existing approaches, particularly in spatial and environmental applications involving censored responses.
@@
The remainder of this document is organized as follows. In Section 2, we provide a review of the SOV and MET methods in a unified framework to set up the mathematical notation for our proposed method. In Section 3, we introduce our transformations of the SOV and the MET integrands to achieve linear complexity for MC sampling. In Section 4, we derive the gradient and Hessian of our transformed MET integrand and show that solving for the proposal density used in MET can be achieved in $\mathcal{O}(n)$ time. In Section 5, we introduce two optional techniques for improving accuracy, namely variable reordering and multi-level MC (Giles, 2008), and we reduce the complexity of the univariate variable reordering from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n^{2})$ through the Vecchia approximation. Section 6 provides numerical comparisons to existing methods for estimating MVN probabilities and sampling TMVN distributions. Section 7 shows that our proposed method enables scalable estimation and inference for partially censored MVN models. Section 8 concludes. All proofs can be found in Appendix LABEL:app:proofs. Our R package, VeccTMVN, which is available on CRAN, implements our methods, while code to reproduce our results can be found at https://github.com/JCatwood/TMVN_Vecchia.
@@ 2 Importance-Sampling Views of SOV and MET
$\Phi_{n}(\mathbf{a},\mathbf{b};\bm{\mu},\bm{\Sigma})$ $=\int_{\mathbf{a}}^{\mathbf{b}}\phi_{n}(\mathbf{x};\bm{\mu},\bm{\Sigma})\mathrm{d}\mathbf{x}$ (1)
$=\int_{\mathbf{a}}^{\mathbf{b}}\frac{1}{(2\pi)^{n/2}|\bm{\Sigma}|^{1/2}}\exp\left(-\frac{1}{2}(\mathbf{x}-\bm{\mu})^{\top}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})\right)\mathrm{d}\mathbf{x},$ (2)
@@
$q(\mathbf{y};\bm{\gamma})={}$ $\prod_{i}^{n}\frac{\mathbbm{1}_{y_{i}\in(\tilde{a}_{i},\tilde{b}_{i})}}{\Phi(\tilde{b}_{i}-\gamma_{i})-\Phi(\tilde{a}_{i}-\gamma_{i})}\phi(y_{i}-\gamma_{i}),$ (9)
$h(\mathbf{y};\bm{\gamma})={}$ $\prod_{i}^{n}\left\{\Phi(\tilde{b}_{i}-\gamma_{i})-\Phi(\tilde{a}_{i}-\gamma_{i})\right\}\frac{\phi(y_{i})}{\phi(y_{i}-\gamma_{i})}\mathbbm{1}_{y_{i}\in(\tilde{a}_{i},\tilde{b}_{i})}.$ (10)
where $\bm{\gamma}$ are the parameters for the proposal density used in MET, and $\phi(\cdot)$ and $\Phi(\cdot)$ are the probability density function (PDF) and the cumulative probability function (CDF) of the univariate standard normal distribution, respectively. From the perspective of importance sampling, (4) and (6) correspond to the SOV and the MET methods, respectively. Specifically, $q(\mathbf{y};\bm{\gamma})$ is the proposal density and $h(\mathbf{y};\bm{\gamma})$ is the density ratio between $\phi_{n}(\mathbf{y};\mathbf{I}_{n})$ truncated in $\mathbf{a}\leq\mathbf{L}\mathbf{y}\leq\mathbf{b}$ and $q(\mathbf{y};\bm{\gamma})$, up to a normalizing constant. MET samples $\mathbf{y}$ from $q(\mathbf{y};\bm{\gamma})$ and then approximates $E_{q(\mathbf{y};\bm{\gamma})}[h(\mathbf{y};\bm{\gamma})]$ by the sample average. Unlike TMVN distributions, $q(\mathbf{y};\bm{\gamma})$ can be sampled as a sequence of $n$ univariate truncated normal distributions as implied by (9). Hence, we can apply accept-reject sampling to TMVN distributions by using $q(\mathbf{y};\bm{\gamma})$ as the proposal density to draw samples from $\phi_{n}(\mathbf{x};\bm{\Sigma})$ truncated in $[\mathbf{a},\mathbf{b}]$. SOV is a special case of MET when the exponential tilting parameters $\bm{\gamma}=\mathbf{0}$. Botev (2017) proposed to use
$\hat{\bm{\gamma}}=\operatorname*{arg\,min}_{\bm{\gamma}}\max_{\mathbf{y}:\,\mathbf{a}\leq\mathbf{L}\mathbf{y}\leq\mathbf{b}}\log h(\mathbf{y};\bm{\gamma}),$ (11)
which minimizes the worst likelihood ratio if we are to draw samples from the TMVN distribution. Solving for $\hat{\bm{\gamma}}$ is discussed in Section 4.
@@ 3.1 Conditional mean and variance under Vecchia
$E[x_{i}|\mathbf{x}_{1:i-1}]=\sum_{j=1}^{i-1}L_{ij}y_{j},\quad\mathrm{Var}[x_{i}|\mathbf{x}_{1:i-1}]=L_{i,i}^{2},$ (12)
where we use the notation $\mathbf{x}_{i:j}=[x_{i},x_{i+1},\ldots,x_{j}]^{\top}$.
@@
where $\eta_{i}=E[x_{i}|\mathbf{x}_{c(i)}]$ and $l_{i}^{2}=\mbox{Var}[x_{i}|\mathbf{x}_{c(i)}]$ can be computed straightforwardly based on an $(m+1)\times(m+1)$ submatrix of $\bm{\Sigma}$. Each conditioning set is typically chosen as the indices of the $m$ nearest previously ordered neighbors. For covariances $\bm{\Sigma}_{i,j}=\mathcal{K}(\mathbf{s}_{i},\mathbf{s}_{j})$ based on isotropic kernels, Euclidean distance between inputs $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ can be used to define nearest neighbors; for other covariance matrices, the correlation distance $(1-|\rho_{ij}|)^{1/2}$ with $\rho_{ij}=\bm{\Sigma}_{ij}/(\bm{\Sigma}_{ii}\bm{\Sigma}_{jj})^{1/2}$ implicitly selects neighbors in a suitable transformed input space (Kang & Katzfuss, 2023). The Vecchia approximation is based on the screening effect (Stein, 2002), which describes the phenomenon that the marginal information gain from far-away responses is often minimal after conditioning on nearby responses. Any Vecchia approximation of a MVN density results in a valid joint MVN density with the same mean but a (slightly) different covariance matrix that has a sparse inverse Cholesky factor. Specifically, Schäfer et al. (2021) showed that, for each $i=1,\ldots,n$, the nonzero entries in the $i$th column of the inverse Cholesky factor $\mathbf{V}=\mathbf{L}^{-\top}$ can be obtained as:
@@
where $\bar{c}(i)=[i,c(i)^{\top}]^{\top}$ and $\mathbf{e}_{1}$ is a vector whose first entry is 1 and all other entries are zero. Hence, $\mathbf{V}$ can be computed using $\mathcal{O}(nm^{3})$ operations and its columns can be computed in parallel. Schäfer et al. (2021) also showed that the covariance matrix under the Vecchia approximation minimizes the KL divergence from the original MVN distribution subject to the implied sparsity constraint on the inverse Cholesky factor $\mathbf{V}$. Furthermore, Vecchia approximations can achieve $\epsilon$-accurate approximations for Matérn-type kernels up to boundary effects using $m$ that only grows polylogarithmically with $n$ (Schäfer et al., 2021; Kang et al., 2024). Numerical evidence that a small $m\leq 50$ is typically sufficient for many spatial and GP settings with very large $n$ is also abundant in the literature (see Section 1 for references). Therefore, in such settings, we can approximately view $m$ as fixed as $n$ grows.
The conditional expectation and variance in Proposition 1 can be computed in $\mathcal{O}(m)$ time for each $i$ when applying the Vecchia approximation to the covariance, resulting in a time complexity of $\mathcal{O}(nm)$ per MC sample.
@@
Proposition 2 is similar to existing results (e.g., Rue & Held, 2005, Thm. 2.3). Its proof is included in the appendix for completeness. The definition of $\mathbf{l}$ in Proposition 2 coincides with that above, namely the univariate conditional standard deviations. With the above expressions for the conditional mean and standard deviation, we can now transform the SOV and MET algorithms such that MC sampling can be performed at linear complexity. Notice that our definitions of $\mathbf{A}$ and $\mathbf{l}$ share similarities with Bolin & Lindgren (2015), which re-parameterized the SOV method with the inverse Cholesky factor, but Bolin & Lindgren (2015) computed $\mathbf{V}$ from the sparse precision matrix, hence leading to higher complexity and less sparsity in $\mathbf{V}$. The conditional expectations and variances computed as proposed in Proposition 2 result in biased MVN probabilities. However, the bias is shown in Proposition 3 to converge to zero exponentially with the number of non-zero entries in $\mathbf{V}$ under certain regularity conditions, including $\bm{\Sigma}$ being based on classes of covariance functions that include Matérn-type kernels up to boundary effects (Kang et al., 2024).

### v3

# JASA-supplement

@@ 1 Introduction
Current computational approaches for MVN probabilities are mostly based on the separation of variable (SOV) method (Genz, 1992). SOV applies a series of transformations such that the MVN probability can be written as an expectation of quantities with low variance, allowing an approximation of the expectation as the average of Monte Carlo (MC) samples. SOV consists of two stages: in the pre-processing stage, the Cholesky factor of the $n\times n$ MVN covariance matrix is computed in $\mathcal{O}(n^{3})$ time; in the MC sampling stage, $N$ samples are drawn, each in $\mathcal{O}(n^{2})$ time. The resulting SOV time complexity is $\mathcal{O}(n^{3}+Nn^{2})$. To improve upon the original SOV method, Gibson et al. (1994) introduced a univariate variable reordering technique used in the pre-processing stage that produces the Cholesky factor and further reduces the MC variance. Based on the series of integrand transformations used in the SOV method, Botev (2017) used importance sampling with minimax exponentially tilted (MET) normal proposal densities during the MC sampling stage to reduce variance. This results in substantially higher accuracy for tail MVN probabilities than SOV, but it requires iterative optimization of the proposal density in the pre-processing stage at a per-iteration complexity of $\mathcal{O}(n^{3})$; in practice, this optimization can be more expensive than the MC sampling for moderately large $n$ (e.g., $n>1{,}000$). Sampling from truncated MVN (TMVN) distributions is closely connected to approximating MVN probabilities (Botev, 2017). Specifically, the integrand in MET (or SOV) can be decomposed into two parts that can be viewed as a proposal density and a density ratio, which can be used for importance sampling from TMVN distributions at the same cost as estimating the corresponding MVN probability. MET is currently the state-of-the-art in terms of achieving the highest accuracy for estimating MVN probabilities and the highest acceptance rate for sampling TMVN distributions.
@@
For evaluating high-dimensional MVN probabilities with spatial covariance matrices, Genton et al. (2018) and Cao et al. (2021) used hierarchical-matrix approximations to reduce the complexity of pre-processing and MC sampling for SOV by a factor of $n^{1/2}$. Bolin & Lindgren (2015) re-parameterized the SOV method with the inverse Cholesky factor to take advantage of the sparse precision matrix under Gaussian Markov random fields. However, they found this approach to often be computationally inefficient, because the sparsity was not well preserved during Cholesky factorization. Other recent approaches for spatial MVN probabilities have used versions of the Vecchia approximation (Vecchia, 1988), which decomposes a multivariate density into a product of univariate conditional densities with small conditioning sets. Saha et al. (2022) applied a Vecchia-type approximation to reduce the complexity of computing the Cholesky factor in the preprocessing stage, but its per-sample complexity remains as $\mathcal{O}(n^{2})$. Nascimento & Shaby (2022) decomposed the $n$-dimensional MVN probability into a product of lower-dimensional MVN probability ratios by assuming conditional independence in the MVN cumulative distribution function (CDF); while this is based on a similar principle as the Vecchia approximation, it appears that the screening effect (i.e., far-away information contributes little after conditioning on nearby points) (Stein, 2002) is weaker for MVN CDFs than for densities, resulting in substantially increased approximation error for CDFs as will be shown empirically in this paper.
@@
Our proposed methods are also based on the Vecchia approximation. A Vecchia approximation of a MVN distribution results in a valid MVN distribution that has a covariance matrix with a sparse inverse Cholesky factor (Datta et al., 2016; Katzfuss & Guinness, 2021). Vecchia approximations can provide asymptotically exact inference in near-linear time for GPs with Matérn kernels subject to boundary effects and smoothness conditions (Schäfer et al., 2021; Kang et al., 2024) — such kernels and their relation to sparse precision matrices were also studied in Lindgren et al. (2011). Numerically, Vecchia approximations have been shown to be highly accurate even with small conditioning-set sizes in many settings, including spatial statistics and high-dimensional and multi-output GPs (e.g., Datta et al., 2016; Katzfuss et al., 2022; Schäfer et al., 2021; Cao et al., 2023). Vecchia approximations can be applied to any covariance matrix, and we expect them to be most accurate if the covariance can be viewed as a stationary kernel with relatively small noise-to-signal ratio defined over some (potentially unknown) input space (Kang & Katzfuss, 2023).
@@
We propose scalable evaluation of MVN probabilities and sampling from TMVN distributions in linear time by applying suitable Vecchia approximations to the MET approach. Hence, we expect our methods to work well in the settings discussed above. We show that the dominant computations for MET and SOV are equivalent to calculating the conditional means and variances of the integration variables. We take full advantage of Vecchia’s sparse inverse Cholesky factor when computing the transformed MET integrand as well as its gradient and Hessian, which allows us to optimize the MET proposal density and sample from TMVN distributions in linear time, an $\mathcal{O}(n^{2})$ reduction compared to MET. We demonstrate theoretically that the Vecchia approximation error decreases exponentially with the number of nonzeros under certain regularity conditions. Numerically, we show that under commonly used spatial kernels, the Vecchia error is often negligible relative to the Monte-Carlo error; for non-decaying kernels, such as constant correlation, the Vecchia approximation still provides a favorable trade-off between accuracy and computation cost. We show that our proposed method facilitates model estimation and posterior inference for partially censored MVN models that offer substantial information gain over existing approaches, particularly in spatial applications involving censored responses.
@@
The remainder of this document is organized as follows. In Section 2, we provide a review of the SOV and MET methods in a unified framework to set up the mathematical notation for our proposed method. In Section 3, we introduce our transformations of the SOV and the MET integrands to achieve linear complexity for MC sampling. In Section 4, we derive the gradient and Hessian of our transformed MET integrand and show that solving for the proposal density used in MET can be achieved in $\mathcal{O}(n)$ time. In Section 5, we introduce two optional techniques for improving accuracy, namely variable reordering and multi-level MC (Giles, 2008). Section 6 provides numerical comparisons to existing methods for estimating MVN probabilities and sampling TMVN distributions. Section 7 shows that our proposed method enables scalable estimation and inference for partially censored MVN models. Section 8 concludes. All proofs, a discussion on multivariate Student-$t$ (MVT) probabilities as well as variable reordering, and additional numerical results an can be found in the appendix. Our R package, VeccTMVN, which is available on CRAN, implements our methods, while code to reproduce our results can be found at https://github.com/JCatwood/TMVN_Vecchia.
@@ 2 Importance-Sampling Views of SOV and MET
$\Phi_{n}(\mathbf{a},\mathbf{b};\bm{\mu},\bm{\Sigma})=\int_{\mathbf{a}}^{\mathbf{b}}\phi_{n}(\mathbf{x};\bm{\mu},\bm{\Sigma})\mathrm{d}\mathbf{x}=\int_{\mathbf{a}}^{\mathbf{b}}\frac{1}{(2\pi)^{n/2}|\bm{\Sigma}|^{1/2}}e^{-\frac{1}{2}(\mathbf{x}-\bm{\mu})^{\top}\bm{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})}\mathrm{d}\mathbf{x},$ (1)
@@
$q^{B}(\mathbf{y};\bm{\gamma})={}$ $\prod_{i}^{n}\frac{\mathbbm{1}_{y_{i}\in(\tilde{a}_{i},\tilde{b}_{i})}}{\Phi(\tilde{b}_{i}-\gamma_{i})-\Phi(\tilde{a}_{i}-\gamma_{i})}\phi(y_{i}-\gamma_{i}),$ (8)
$h^{B}(\mathbf{y};\bm{\gamma})={}$ $\prod_{i}^{n}\left\{\Phi(\tilde{b}_{i}-\gamma_{i})-\Phi(\tilde{a}_{i}-\gamma_{i})\right\}\frac{\phi(y_{i})}{\phi(y_{i}-\gamma_{i})}\mathbbm{1}_{y_{i}\in(\tilde{a}_{i},\tilde{b}_{i})}.$ (9)
where $\bm{\gamma}$ are the parameters for the proposal density used in MET, and $\phi(\cdot)$ and $\Phi(\cdot)$ are the probability density function (PDF) and the cumulative probability function (CDF) of the univariate standard normal distribution, respectively. From the perspective of importance sampling, (3) and (5) correspond to the SOV and the MET methods, respectively. Specifically, $q^{B}(\mathbf{y};\bm{\gamma})$ is the proposal density and $h^{B}(\mathbf{y};\bm{\gamma})$ is the density ratio between $\phi_{n}(\mathbf{y};\mathbf{I}_{n})$ truncated in $\mathbf{a}\leq\mathbf{L}\mathbf{y}\leq\mathbf{b}$ and $q^{B}(\mathbf{y};\bm{\gamma})$, up to a normalizing constant. MET samples $\mathbf{y}$ from $q^{B}(\mathbf{y};\bm{\gamma})$ and then approximates $E_{q^{B}(\mathbf{y};\bm{\gamma})}[h^{B}(\mathbf{y};\bm{\gamma})]$ by the sample average. Unlike TMVN distributions, $q^{B}(\mathbf{y};\bm{\gamma})$ can be sampled as a sequence of $n$ univariate truncated normal distributions as implied by (8). Hence, we can apply accept-reject sampling to TMVN distributions by using $q^{B}(\mathbf{y};\bm{\gamma})$ as the proposal density to draw samples from $\phi_{n}(\mathbf{x};\bm{\Sigma})$ truncated in $[\mathbf{a},\mathbf{b}]$. SOV is a special case of MET when the exponential tilting parameters $\bm{\gamma}=\mathbf{0}$. Using
$\hat{\bm{\gamma}}=\operatorname*{arg\,min}_{\bm{\gamma}}\max_{\mathbf{y}:\,\mathbf{a}\leq\mathbf{L}\mathbf{y}\leq\mathbf{b}}\log h^{B}(\mathbf{y};\bm{\gamma}),$ (10)
was proposed by Botev (2017) to minimize the worst likelihood ratio if we are to draw samples from the TMVN distribution. Solving for $\hat{\bm{\gamma}}$ is discussed in Section 4.
@@ 3.1 Conditional mean and variance under Vecchia
$E[x_{i}|\mathbf{x}_{1:i-1}]=\sum_{j=1}^{i-1}L_{ij}y_{j},\quad\mathrm{Var}[x_{i}|\mathbf{x}_{1:i-1}]=L_{i,i}^{2},\ \mbox{where }\mathbf{x}_{i:j}=[x_{i},x_{i+1},\ldots,x_{j}]^{\top}.$ (11)
@@
where $\eta_{i}=E[x_{i}|\mathbf{x}_{c(i)}]$ and $l_{i}^{2}=\mbox{Var}[x_{i}|\mathbf{x}_{c(i)}]$ can be computed straightforwardly based on an $(m+1)\times(m+1)$ submatrix of $\bm{\Sigma}$. Each conditioning set is typically chosen as the indices of the $m$ nearest previously ordered neighbors. For covariances $\bm{\Sigma}_{i,j}=\mathcal{K}(\mathbf{s}_{i},\mathbf{s}_{j})$ based on isotropic kernels, Euclidean distance between inputs $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ can be used to define nearest neighbors; for other covariance matrices, the correlation distance $(1-|\rho_{ij}|)^{1/2}$ with $\rho_{ij}=\bm{\Sigma}_{ij}/(\bm{\Sigma}_{ii}\bm{\Sigma}_{jj})^{1/2}$ implicitly selects neighbors in a suitable transformed input space (Kang & Katzfuss, 2023). The Vecchia approximation is based on the screening effect (Stein, 2002), which describes the phenomenon that the marginal information gain from far-away responses is often minimal after conditioning on nearby responses. Schäfer et al. (2021) showed that, for each $i=1,\ldots,n$, the nonzero entries in the $i$th column of the inverse Cholesky factor $\mathbf{V}=\mathbf{L}^{-\top}$ can be obtained as:
@@
where $\bar{c}(i)=[i,c(i)^{\top}]^{\top}$ and $\mathbf{e}_{1}$ is a vector whose first entry is 1 and all other entries are zero. Hence, $\mathbf{V}$ can be computed using $\mathcal{O}(nm^{3})$ operations and in parallel. Numerical evidence that a small $m\leq 50$ is typically sufficient for many spatial and GP settings with very large $n$ is also abundant in the literature (see Section 1 for references). Therefore, in such settings, we can approximately view $m$ as fixed as $n$ grows. The conditional expectation and variance in Proposition 1 can be computed in $\mathcal{O}(m)$ time for each $i$ when applying the Vecchia approximation to the covariance, resulting in a per-sample complexity of $\mathcal{O}(nm)$.
@@
Proposition 2 is similar to existing results (e.g., Rue & Held, 2005, Thm. 2.3). Its proof is included in the appendix for completeness. The definition of $\mathbf{l}$ in Proposition 2 coincides with that above, namely the univariate conditional standard deviations. Notice that our definitions of $\mathbf{A}$ and $\mathbf{l}$ share similarities with Bolin & Lindgren (2015), which re-parameterized the SOV method with the inverse Cholesky factor, but Bolin & Lindgren (2015) computed $\mathbf{V}$ from the sparse precision matrix, hence leading to higher complexity and less sparsity in $\mathbf{V}$. The conditional expectations and variances computed as proposed in Proposition 2 result in biased MVN probabilities. However, the bias is shown in Proposition 3 to decrease exponentially with the number of non-zero entries in $\mathbf{V}$ under certain regularity conditions, covering Matérn kernels with smoothness $\nu$ such that $\nu+d/2$ is an integer, where $d$ is the dimension of the domain over which Matérn kernels are defined (Schäfer et al., 2021).